OpenStudy (anonymous):

In the xy-plane, a particle moves along the parabola y=x^2-x with a constant speed of 2(10)^1/2 units per second. If dx/dt>0, what is the value of dy/dt when the particle is at the point (2,2)?

OpenStudy (anonymous):

The choices I was given are a) 2/3 b) 2(10^.5)/3 c) 6

OpenStudy (dumbcow):

\[v^{2} = (\frac{dx}{dt})^{2}+(\frac{dy}{dt})^{2}\] \[\frac{dx}{dt} = \frac{dx}{dy}*\frac{dy}{dt}\] \[\frac{dx}{dy} = \frac{1}{2x-1}\] ...what is the speed? 2*sqrt10 , i am not getting any of your answers

OpenStudy (anonymous):

Yeah, that's it...It's from one of the AP Calculus BC Released Exams.

OpenStudy (dumbcow):

anyway the equation then becomes \[v = \frac{dy}{dt}\sqrt{1+\frac{1}{(2x-1)^{2}}}\] v = 2*sqrt10 solve for dy/dt

OpenStudy (anonymous):

Thank you!!

OpenStudy (dumbcow):

your welcome..ahh i see what i did wrong...answer is 6

OpenStudy (anonymous):

cool. just plugged it in & i got that too.